OpenStudy (anonymous):

Write the expression as a single logarithm. 5log3u+8log3v= Simplify answer

OpenStudy (hoblos):

= log(3u)^5 + log(3v)^8 =log[(3u)^5 * (3v)^8]

OpenStudy (hoblos):

welcome :)

myininaya (myininaya):

is 3 the base? \[5\log_3(u)+8\log_3(v)=\log_3(u^5)+\log_3(v^8)=\log_3(u^5 \cdot v^8)\]

OpenStudy (anonymous):

yes it is